OpenStudy (anonymous):

evaluate: log2 16 + log2 64

OpenStudy (dumbcow):

use log property: log a + log b = log ab

OpenStudy (saifoo.khan):

Right.

OpenStudy (dumbcow):

actually in this case, it would be better to rewrite 16 and 64 as powers of 2

OpenStudy (saifoo.khan):

therwise my calculate will work on that.

OpenStudy (anonymous):

then use change of base rule

OpenStudy (saifoo.khan):

otherwise*

OpenStudy (anonymous):

or just say that \[\log_2(16)=4\] since \[2^4=16\] and \[\log_2(64)=6\] since \[2^6=64\] and \[4+6=10\]

OpenStudy (anonymous):

change of base???!!!

OpenStudy (anonymous):

loga b=logc b/logca